OpenStudy (anonymous):

Simplify the expression. –12 • 1/6 • 3 + 48 ÷ 3 A. –10 B. 10 C. 14 D. –34

OpenStudy (anonymous):

@AloneS.

OpenStudy (anonymous):

yeah she never has talked to me ive tryed to add her to my questions and tryed to start a conversation with her but never worked so but you seem nice so yeah

OpenStudy (anonymous):

sorry for the bad grammar

OpenStudy (alones.):

np

OpenStudy (anonymous):

do you know how to help with this question?

OpenStudy (alones.):

-12 * 1/6 = -2

OpenStudy (anonymous):

what is the 1/6 is it a fraction?

OpenStudy (alones.):

-2 * 3 = -6

OpenStudy (alones.):

yes, 1/6 is a fraction

OpenStudy (anonymous):

oh ok they put it like that in my question

OpenStudy (anonymous):

so i was confused

OpenStudy (alones.):

yeah... -6 + 48 = 42

OpenStudy (anonymous):

wait PEMDOS

OpenStudy (anonymous):

PEMDAS

OpenStudy (alones.):

42 divided by 3 = 14 Yes, PEMDAS. -12 * 6 is first

OpenStudy (alones.):

So, the answer is 14. I'm pretty sure.

OpenStudy (anonymous):

oh thank you i have a lot more questions so can you help?

OpenStudy (alones.):

I can try, yeah.

OpenStudy (anonymous):

k ill post then tag you k

OpenStudy (alones.):

I'm going t switch account though. I'm going to be in wolfman77 account, okay?

OpenStudy (anonymous):

k